### 15 Feb 2024 v1

# CodeMind: A Framework to Challenge Large Language Models for Code Reasoning

Abstract
Solely relying on test passing to evaluate Large Language Models (LLMs) for code synthesis may result in unfair assessment or promoting models with data leakage. As an alternative, we introduce CodeMind, a framework designed to gauge the code reasoning abilities of LLMs. CodeMind currently supports three code reasoning tasks: Independent Execution Reasoning (IER), Dependent Execution Reasoning (DER), and Specification Reasoning (SR). The first two evaluate models to predict the execution output of an arbitrary code or code the model could correctly synthesize. The third one evaluates the extent to which LLMs implement the specified expected behavior.
Our extensive evaluation of nine LLMs across five benchmarks in two different programming languages using CodeMind shows that LLMs fairly understand control flow constructs and, in general, are capable of reasoning how inputs evolve to output, specifically for simple programs and the ones they can correctly synthesize. However, their performance drops for code with higher complexity, non-trivial logical and arithmetic operators, non-primitive types, and API calls. Furthermore, we observe that, while correlated, specification reasoning (essential for code synthesis) does not imply execution reasoning (essential for broader programming tasks such as testing and debugging): ranking LLMs based on test passing can be different compared to code reasoning.
Department of Computer Science
University of Illinois at Urbana-Champaign, Illinois, US
1 Introduction
Large Language Models (LLMs) have shown exceptional programming abilities, specifically when instruction-tuned or prompted through Chain- or Tree-of-Thoughts (CoT (Wei et al., 2022b) or ToT (Yao et al., 2023)) and in-context learning (Wei et al., 2022a; Garg et al., 2022). However, several studies suggest that LLMs struggle to generalize this exceptional ability, specifically when the dataset becomes more complex (Du et al., 2023; Jimenez et al., 2023), or the task requires understanding code, rather than natural language (Pan et al., 2023; Min et al., 2023). This is mainly because LLMs are trained to associate code synthesis with natural language specifications, i.e., reason how to combine code constructs similar to examples they have seen while satisfying requirements explained in the specification.
To illustrate how code reasoning tasks can evaluate LLMs, Figure 1-a shows a code synthesized by GPT-3.5 given natural language specification. The code constructs corresponding to the specification are highlighted with matching colors. Due to the ambiguity in the natural language, this code returns the smallest number in the list rather than the number at the index equal to the value of the smallest number. As a result, for a given input $[2,5,4,3]$, the code returns 2 instead of 4, and the assertion fails.
Figure 1: An example illustrating the importance of evaluating LLMs on code reasoning
One way to assess the inductive code reasoning of LLMs is to include specific expected program behavior and check whether the generated code can reproduce that behavior. This entails a level of code reasoning, which we refer to as Specification Reasoning (SR). Figure 1-b shows the new specification and the corresponding generated code. Executing the code given the specified input-output pair results in a test pass, indicating the ability of GPT-3.5 to understand the given specification and generate a correct code.
Including test data in prompts has been a known practice to improve the performance of models in programming tasks (Chen et al., 2022; Zhong et al., 2022; Shi et al., 2022; Zhang et al., 2023). However, it is a weak proxy for code reasoning as it still involves the association of code and natural language. A deeper level of code reasoning is reasoning about execution output given an input, which we call Execution Reasoning (ER). This task challenges LLMs more, requiring them to reason about code without any natural language cross reference. Figure 1-c shows the CoT reasoning of GPT-3.5 in response to the ER task. Even though the model could generate a code that produced the expected output (and is correct if validated through testing), it cannot correctly reason about the code execution given the same inputs to predict the output.
To automate code reasoning assessment, we propose CodeMind. CodeMind currently offers three inductive code reasoning tasks: Independent Execution Reasoning (IER) and Dependent Execution Reasoning (DER) assess if LLMs can reason about how given inputs evolve to output for any arbitrary code (IER) or only the code that it correctly synthesized. Specification Reasoning (SR) evaluates the extent to which LLMs can reason and implement the specified behavior.
Using CodeMind, we performed a large-scale ground theory study to assess LLMs for code reasoning. We selected nine models, including both general-purpose and Code LLMs and prompted them for IER, DER, and SR tasks on 5395 programs written in Java and Python [^1] These programs are from five programming benchmarks, namely HumanEval (Chen et al., 2021), MBPP (Odena et al., 2021), CRUXEval (Gu et al., 2024) CodeNet (Puri et al., 2021), and Avatar (Ahmad et al., 2021).. We observe that:
(1) LLMs have a good grasp of code constructs, likely due to alignment with concepts in the natural language specification. The instruction-tuned models can explain the code statement by statements and follow the execution of the programs in general. LLM code reasoning abilities, however, are limited to simple programs. Furthermore, models such as GPT-3.5 and MagicCoder (Wei et al., 2023), although they correctly explain what the code does, may fail to keep track of data flow and correctly reason about execution output. Open-source LLMs that have achieved comparable effectiveness as GPT models in code synthesis (Wei et al., 2023; Roziere et al., 2023; Luo et al., 2023) are behind them with a huge gap concerning code reasoning (section 5).
(2) LLMs can reason about test data in the specification, even if deceptive, and bring that into the reasoning process for code synthesis (section 7). However, their reasoning is bottlenecked by their inherent limitation. They achieve a higher performance reasoning about the code they can correctly synthesize (section 6).
(3) On a dataset with complex programs, there is a negligible to no correlation between the ranking of models based on code synthesis—generating a code that passes all tests—and code reasoning performance (section 6). This necessitates CodeMind tasks and metrics to complement the evaluation of LLMs for code.
(4) Nested code constructs, complex conditional predicates and loop conditions, non-trivial arithmetic and logic operators, and API invocations can significantly challenge LLMs for code reasoning (section 8).
Our contributions are: (1) CodeMind framework for code reasoning that formally defines three inductive code reasoning tasks. CodeMind is open-source (CodeMind, 2024) and accepts contributions from researchers to integrate more code reasoning tasks into it. (2) A large-scale ground-theory evaluation of LLMs for code reasoning using CodeMind. (3) A comprehensive, in-depth analysis of results that offers a catalog of root causes negatively impacting the abilities of LLMs for code reasoning. This catalog would be a valuable guideline for developing better benchmarks that truly evaluate the programming abilities of LLMs.
2 Related Work
A large body of work has assessed LLMs for reasoning tasks of different modalieties (Deshpande et al., 2021; Wu et al., 2023; Miceli-Barone et al., 2023; Bubeck et al., 2023; Wang et al., 2023; Imani et al., 2023; Luo et al., 2023; Huang et al., 2023; Valmeekam et al., 2022; Min et al., 2023), including natural language, visual data, math, logic, and code. CodeMind is more closely related to the very recent studies focusing on code reasoning (La Malfa et al., 2024; Gu et al., 2024; Zhang et al., 2024).
CRUXEval benchmark is a concurrent work investigating the problem of code reasoning abilities of LLMs using a dataset of simple programs generated by CodeLlama (34B) with test cases  (Gu et al., 2024). They evaluated a series of LLMs on CRUXEval for input and output prediction tasks. Compared to CRUXEval, CodeMind proposes more inductive code reasoning tasks, includes more programs with a variety of levels of complexity, and controls between code synthesis and reasoning tasks by evaluating LLMs using the same program. CodeMind is also equipped with a static analysis pipeline to enable in-depth examination and drawing informed conclusions.
La Malfa et al. (2024) evaluate LMs to predict variable values at each code statement. Our experiments are larger compared to them: more programs with a diverse distribution of complexity and different programming languages, and more studied LLMs. We also offer more code reasoning tasks and present a cross-analysis of code synthesis and reasoning abilities.
Zhang et al. (2024) investigate transformers’ ability to learn or infer the recursive patterns from input and output pairs. They conclude that due to the inherent limitations of transformers, they may fail to learn recursion and instead find shortcut algorithms to reason about how outputs are related to inputs. Compared to this work, we evaluate LLMs regardless of architecture and training data but from the program perspective. We show LLMs can understand recursion but usually lose track of data flow due to the inability to correctly reason about loop conditions.
3 CodeMind
Program specification defines a function $S:S_{I}\rightarrow S_{O}$, where $S_{I}$ is a set of all possible inputs to the program and $S_{O}$ is a set of corresponding outputs. A code synthesized based on the implementation is usually a (partial [^2] Due to the specification’s abstraction and ambiguities, especially when given in natural language, implementation is a partial function of what specification defines.) function $C:C_{I}\rightarrow C_{O}$. We define a program to be correct with respect to specification, if it satisfies all the following conditions:
$C_{I}\subseteq S_{I}$, $C_{O}\subseteq S_{O}$, $\forall i\in C_{I},C(i)=S(i)$
This entails the models to reason about how inputs evolve to a given output through implementation (execution reasoning) and implements a code such that it generates correct outputs for a given input (specification reasoning).
3.1 Execution Reasoning
Considering the aforementioned formalization, we define two execution reasoning tasks as follows.
Definition 1: Independent Execution Reasoning (IER). Given a program $C:C_{I}\rightarrow C_{O}$ and set of inputs $\hat{I}=\{i|i\in C_{I}\}$, LLM $L$ can correctly reason about code execution if $\hat{o}=C(\hat{I})$, where $\hat{o}=L(\hat{I})$ is the predicted output by $L$. Note that in this task, we do not deal with specification, so we can assess LLMs for any arbitrary code that we have ground-truth pairs of $\langle\hat{I},\hat{o}\rangle$.
IER evaluates LLMs for any arbitrary code for general inductive code reasoning, which requires understanding code constructs, arithmetic and logic operations, and control flow. However, even for human developers, reasoning about their developed code is easier than any arbitrary code. Furthermore, as a self-consistency (Min et al., 2023) measurement, LLMs should be able to reason about the code they can correctly synthesize. This demands to have the following execution reasoning task.
Definition 2: Dependent Execution Reasoning (DER). Given a specification $S:S_{I}\rightarrow S_{O}$, a program $C:C_{I}\rightarrow C_{O}$ generated by LLM $L$, and set of inputs $\hat{I}=\{i|i\in C_{I},C(i)=S(i)\}$, LLM $L$ can correctly reason about code execution if $\hat{o}=C(\hat{I})$, where $\hat{o}=L(\hat{I})$ is the predicted output by $L$. The assumption here is that when LLM $L$ generates code $C$ that passes the test $\langle\hat{I},\hat{o}\rangle$, it be able to predict $\hat{o}$ correctly.
3.2 Specification Reasoning
In addition to inductive execution reasoning, a model should understand specification to synthesize a correct code. We formally define the specification reasoning task as follows.
Definition 3: Specification Reasoning (SR). Given a specification $S:S_{I}\rightarrow S_{O}$, an arbitrary $\langle i,o\rangle$ specified in the prompt along with the natural language specification, where $i\in S_{I},o\in S_{O},S(i)=o$, and program $C:C_{I}\rightarrow C_{O}$ generated by LLM $L$, the LLM can correctly reason about specification if $C(i)=S(i)$. In other words, LLM $L$ should be able to pass a test with $\langle i,o\rangle$, when they are explicitly specified in the prompt.
[…]
Dataset Programming Language # Subjects General LLMs Code LLMs
GPT-4 GPT-3.5 Llama 2 Mistral CodeLlama DeepSeekCoder MagicCoder StarCoder WizardCoder
MBPP Python 408 80.88% 71.32% 45.59% 31.37% 42.40% 57.84% 59.80% 43.63% 46.08%
HumanEval Python 162 79.01% 64.20% 30.86% 32.72% 45.06% 41.98% 52.47% 38.89% 40.12%
CruxEval Python 800 80.50% 65.13% 25.38% 34.13% 37.75% 44.38% 46.50% 35.50% 35.88%
CodeNet Python 1914 70.43% 49.06% 18.97% 17.35% 27.95% 26.65% 33.28% 26.28% 24.87%
Java 1939 71.17% 51.93% 23.99% 18.15% 28.52% 32.13% 36.46% 29.34% 29.35%
Avatar Python 86 52.33% 39.53% 24.42% 16.28% 23.26% 18.60% 24.42% 19.77% 24.42%
Java 86 48.84% 34.88% 23.26% 11.63% 27.91% 23.26% 24.42% 13.95% 13.95%
Total Java and Python 5395 72.60% 54.24% 24.26% 21.54% 30.40% 33.85% 38.68% 30.14% 29.99%
3.3 Evaluating Code Reasoning
We measure the performance of models in code reasoning for a given code with the Correct Reasoning Score (CRS), which is $1$ if the model can correctly reason about the code and 00 otherwise. We also introduce the Correct Reasoning Rate (CRR) metric, a collective metric that measures how much a given LLM can reason about multiple programs in a benchmark. We calculate CRR for a set of $m$ programs in benchmark $P$ as:
$CRR(P)=\dfrac{\sum\limits_{i=1}^{m}\llbracket CRS(p_{i}\in P)=1\rrbracket}{m}$
4 Experimental Setup
Our study includes nine LLMs and $5395$ programs in Java and Python programming languages from five programming datasets. We explain the details of LLMs and program selection below.
Subject LLMs. We chose nine pre-trained or instruction-tuned models, covering both general-purpose and Code LLMs. Our choice was limited to computing resources, so we selected models with less than $20$B parameters that outperform the rest for programming tasks. Our subject LLMs are GPT-4 (OpenAI, 2023b), GPT-3.5 (OpenAI, 2023a), Llama 2 (13B) (Touvron et al., 2023), Mistral (Jiang et al., 2023), CodeLlama (13B, instruction-tuned)(Roziere et al., 2023), StarCoder (15.5B)(Li et al., 2023), WizardCoder (15B, instruction-tuned)(Xu et al., 2023), MagicCoder (7B)(Wei et al., 2023) (instruction-tuned), DeepSeekCoder (6.7B)(Bi et al., 2024). We followed the best practices and customized the prompt templates per each model (all prompts are publicly available for further investigation (CodeMind, 2024)). Except for the GPT models, we set the temperature to zero to ensure the reproducibility of the results. Our code is open-source to users for using CodeMind for other models and temperatures.
Subject Programs. Our criteria for selecting subject programs were the existence of test data (inputs and corresponding expected output) and implementations of the same program in multiple programming languages (to investigate its impact on code reasoning). From several existing benchmarks (Wang et al., 2022; Athiwaratkun et al., 2022; Chen et al., 2021; Liu et al., 2023; Gu et al., 2024; Zheng et al., 2023; Cassano et al., 2022; Jimenez et al., 2023; Du et al., 2023; Odena et al., 2021; Puri et al., 2021; Ahmad et al., 2021), we chose the programs in HumanEval (Chen et al., 2021), MBPP (Odena et al., 2021), CodeNet (Puri et al., 2021), Avatar (Ahmad et al., 2021), and CruxEval (Gu et al., 2024). We chose Java and Python versions of the programs as they are more prominently used programming languages. HumanEval and MBPP are well-known benchmarks for code synthesis. CodeNet and Avatar are code translation benchmarks. CRUXEval is a benchmark of relatively simple Python programs generated by CodeLlama (34B) to evaluate input prediction and output prediction of LLMs.
Figure 2: Complexity distribution of the subject programs in terms of Cyclomatic Complexity (CC) and Line of Code (LoC)
Figure 2 shows the complexity distribution of the programs in terms of Cyclomatic Complexity, $CC$ (Gill & Kemerer, 1991), and Lines of Code (LoC). $CC$ measures the number of independent execution paths in the program control flow graph (CFG). The metric is computed for a class as $CC=E-N+2P$, where $E$ and $N$ are the number of edges and nodes in the CFG, respectively, and $P$ is the number of methods in the class. In general, a higher $CC$ indicates a more complex program.
For code reasoning tasks, the model should reason which execution path to take for a given input to predict the output. So, the higher number of independent paths makes it unlikely for the model to succeed by chance. $CC$ might be correlated with the number of lines in the program, but more lines do not cause higher $CC$. For example, a program with $10$ lines and no conditional or loop constructs only has one execution path, while a program with $8$ lines and two nested conditional statements has $3$ or $4$ execution paths, depending on the conditional predicates.
5 LLM Evaluation on IER
Figure 3: Impact of CC on CRR performance of LLMs in IER
To evaluate the performance of LLMs on the IER task, we promoted them under two settings: direct answering and CoT. For direct answering, we prompted each model to predict the output of given inputs. Under the CoT setup, we first instruct the models to simulate the execution step by step by predicting the output value after execution of each statement. We then ask the model to predict the output of given inputs. In both settings, the prompt contains one in-context example for two purposes: introducing the IER task and instructing the response formatting.
Given that IER only requires an arbitrary code and corresponding ground-truth pair of $\langle\hat{I},\hat{o}\rangle$ (section 3.1), we prompted the LLMs using all $5395$ subject programs in this experiment. Table 1 shows the result of this experiment through CoT prompting. From these results, we observe that:
[…]
Dataset # Subjects Task General LLMs Code LLMs
GPT-4 GPT-3.5 Mistral CodeLlama DeepSeekCoder MagicCoder StarCoder WizardCoder
MBPP 408 Synthesis $86.52\%$ $80.39\%$ 43.36% 56.86% 72.30% 70.34% 44.85% 61.03%
Reasoning 82.62% 79.20% 43.50% 43.53% 63.39% 69.34% 56.83% 48.19%
CRR Improvement cf. IER $1.74\%\uparrow$ $7.88\%\uparrow$ $11.89\%\uparrow$ $1.13\%\uparrow$ $5.15\%\uparrow$ $9.54\%\uparrow$ $13.20\%\uparrow$ $2.11\%\uparrow$
HumanEval 162 Synthesis 87.65% 69.75% 52.47% 67.90% 81.48% 79.62% 48.15% 72.46%
Reasoning 80.28% 74.63% 34.12% 35.45% 54.55% 53.49% 58.97% 59.50%
CRR Improvement cf. IER $1.27\%\uparrow$ $10.70\%\uparrow$ $1.4\%\uparrow$ $9.61\%\downarrow$ $12.57\%\uparrow$ $1.02\%\uparrow$ $20.08\%\uparrow$ $19.38\%\uparrow$
  • GPT models outperform others on the IER task, with large margins of $33.92\%$ (GPT-4) and $15.56$ (GPT-3.5) from the best open-source model. Among the open-source models, except for the Avatar dataset, MagicCoder outperforms others with an average margin of $4.83\%$.
  • On the datasets with samples in both Java and Python, all the models experience a performance drop (average drop of $2.91\%$ in CodeNet and $2.33\%$ in Avatar). This is likely because Java enforces a stricter syntax and typing system than Python, making the code execution reasoning harder.
  • Compared to direct answering, CoT prompting, under which the models articulate the execution process verbally before predicting the output, results in $5.24\%$ improvement in the IER performance of the models on average. However, the less-than-ideal accuracy of (open-source) models, even with CoT prompting, demands a fundamental change.
  • Moving down in the table, the models face greater challenges in IER, i.e., reasoning about the execution on CodeNet and Avatar programs, compared to MBPP, HumanEval, and CRUXEval. One potential reason is the complexity of such programs as demonstrated in Figure 2. A detailed breakdown of the model’s performance (Figure 3) shows a strong negative correlation [^3] Spearman’s Rank Order Correlation (ROC) (Spearman, 1961) between CC and CRR. between CRR and CC, confirming that models struggle more in IER for a more complex code. At the same time, some models, namely Llama 2, CodeLlama, MagicCoder, StarCoder, and WizardCoder, achieve a lower performance on CRUXEval compared to HumanEval, which are less complex regarding both LoC and CC. This entails a further better understanding of what factors other than CC impact the CRR performance of the models (section 8).
6 LLM Evaluation on DER
We seek to address the critical question of how effectively the model can correctly reason about the correct programs it has generated. This evaluation requires us to align code synthesis and code reasoning tasks together. Our pipeline for evaluating DER consists of three steps: (1) following the best practices, we prompted subject LLMs for code synthesis; (2) we ran the synthesized program against existing tests; and (3) for the programs with test pass, we prompted the model for code execution reasoning using the chosen test input and under CoT style. Note that we also removed the comments from the synthesized code for fairness. We excluded the programs from CRUXEval, CodeNet, and Avatar, since these datasets are not designed for code synthesis and lack proper program specifications. Also, we could not reproduce the code synthesis results of Llama 2 and excluded that from subject LLMs. Similar to IER experiments, we set the temperature to zero  [^4] As a result of this design decision, our synthesis results might be different from existing leaderboards. to account for the non-determinism and reproducibility of the results.
Table 2 shows the results of this experiment. GPT models still outperform open-source models on the DER task, with a margin of $17.97$ (GPT-4) and $13.13$ (GPT-3.5) from the best open-source model. Compared to IER, the gap between GPT models and open-source models has been reduced. We can also observe that the models achieve $6.84\%$ higher CRR on average in the DER task (except CodeLlama on HumanEval), compared to IER.
Before concluding the models are more competent in the execution reasoning when evaluated on the programs they correctly synthesize, we compared the programs in this experiment with those in IER experiment. If true, lower complexity might be the root cause of higher CRR on the DER task. Figure 4 shows the CC distribution of the programs in MBPP and HumanEval, compared to that generated by subject LLMs. We can observe that the synthesized code, if not more complex, is no less than the ground-truth programs in these datasets. Consequently, we confirm that models reason better on a code they correctly synthesize. However, there is still a considerable gap between the code synthesis and reasoning abilities of the LLM, specifically on open-source models.
Figure 4: CC distribution of the programs synthesized by LLMs compared to the original programs in the HumanEval (top) and MBPP (bottom) datasets
Given that code synthesis and reasoning are unified in DER, we first computed the Spearman’s ROC between the rank of models based on numbers in the Synthesis row and Reasoning row for each dataset. The results show a strong positive correlation on MBPP ($\rho=0.85$), but a negligible correlation on HumanEval ($\rho=0.17$). These results communicate a strong message: the ranking of LLMs based on their code synthesis abilities (pass@k) could be significantly different than their reasoning abilities on the same code. This necessitates a framework such as CodeMind that promotes other evaluation aspects of LLMs for code.
Table 3: Performance of LLMs on SR task. Symbol $\downarrow$ indicates a drop from the previous setting (row above), and $\uparrow$ indicates an increase from the previous setting (row above).
Dataset Setting General LLMs Code LLMs
GPT-4 GPT-3.5 Mistral CodeLlama DeepSeekCoder MagicCoder StarCoder WizardCoder
MBPP With Test 90.69% 85.05% 50.74% 63.73% 78.68% 75.25% 51.47% 67.89%
No Test 72.13% $\downarrow$ 78.87% $\downarrow$ 48.28% $\downarrow$ 53.68% $\downarrow$ 67.65% $\downarrow$ 69.61% $\downarrow$ 41.67% $\downarrow$ 52.21% $\downarrow$
Misleading Test 68.14% $\downarrow$ 74.02% $\downarrow$ 50.74% $\uparrow$ 59.07% $\uparrow$ 68.63% $\uparrow$ 67.40% $\downarrow$ 40.20% $\downarrow$ 58.09% $\uparrow$
HumanEval With Test 91.98% 74.07% 57.41% 70.37% 87.04% 81.48% 56.17% 76.54%
No Test 88.27% $\downarrow$ 70.37% $\downarrow$ 54.32% $\downarrow$ 65.43% $\downarrow$ 82.10% $\downarrow$ 80.86% $\downarrow$ 38.89% $\downarrow$ 76.54%
Misleading Test 83.95% $\downarrow$ 65.43% $\downarrow$ 53.70% $\downarrow$ 61.73% $\downarrow$ 79.63% $\downarrow$ 74.69% $\downarrow$ 27.04% $\downarrow$ 66.05% $\downarrow$
7 Evaluation on SR
Specification Reasoning (SR) offers a novel perspective in understanding the code synthesis process of LLMs, particularly in how they leverage input-output specifications. To evaluate the abilities of LLMs for SR, we prompt LLMs for code synthesis under the following three settings:
Figure 5: CRR of top five best-performing LLMs per specific code constructs across all datasets. We abbreviate the tags with B (Basic), F (For), I (If), NI (Nested If), NL (Nested Loop), S (Switch), T (Try), and W (While)
(1) Natural language specification with one ground-truth input-output. Under this setting, we randomly select one of the existing tests and add that to the specification. We validate the synthesized code using only this test.
(2) Natural language specification with no input-output. We remove the test added to the specification in the previous setting and re-prompt LLMs for code synthesis. We validate the synthesized code using only the test from the previous setting. Intuitively, if including test data can help LLMs in code synthesis, we observe a drop in LLMs’ performance.
(3) Natural language specification with misleading input-output. We mutate the expected output of the test from the first setting and add it to the specification. We validate the synthesized code using the original test. The mutation changes the expected output to a value that does not align with the specification. For example, if the expected output is True, mutation changes it to False. Similarly, if the expected output is a positive integer, we mutate it to a negative one with a large difference. Intuitively, due to the divergence with natural language specification, misleading input-output should further drop LLMs’ performance.
We followed a similar setup for this experiment as the one in section 6; performed this experiment only on MBPP and HumanEval programs. We also pre-processed the prompts from HumanEval, which initially contained input-output samples. The results in Table 3 show that the performance of LLMs in code synthesis is, on average, $7.36\%$ higher with test data included in the specification. Introducing deceptive tests in the specification detrimentally affects the LLMs’ performance in code synthesis compared to a legitimate test ($10\%$ performance drop on average). However, compared to No Test cases, the performance drop across all the models and programs is only $2.65\%$ on average. Regardless, these results showcase the ability of LLMs to reason and utilize the test data in the specification.
Figure 6: Impact of loop length in Java programs (CodeNet and Avatar) on LLMs’ performances
8 In-Depth Analysis of Results
We further analyzed the IER results, which evaluate the general ability of LLMs in code reasoning. In the first step, we wanted to see if LLMs understand different code constructs. Without knowing the logic of each code construct, reasoning about code execution is impossible. To that end, we tagged each of $5395$ programs based on code constructs used in their implementation with the following labels: For, While, If, Try, Switch, Nested Loop, Nested If, and Basic. A program tagged with a Basic label has no special code construct. Next, we clustered the programs per tag and computed the CRR of LLMs for each cluster. Figure 5 shows the results of this analysis for the top five best-performing LLMs. We can observe that models handle conditional statements better than recursion, except for Try-Catch or Try-Except statements. Furthermore, when it comes to nested constructs, the CRR values notably drop. Impact of Loop Properties. Given that models struggle the most with recurring constructs, we focused the programs with For, While, and Nested Loop tags at the next step. We hypothesize this struggle is due to the loop’s length or determining the loop’s length. The former questions whether it is harder for the model to track the program’s data flow as loops get longer. The latter questions the abilities of models to reason about how many times a code block should be repeated, regardless of how long the length would be.
Figure 6 plots the distribution of correct versus incorrect cases and CRR values per each loop length in Java programs. Sub-figure labels show the ROC coefficients between the loop length and CRR. We can observe a moderate to strong negative correlation between the loop length and CRR of the models, i.e., CRR decreases as the loop length increases. By manually investigating the incorrect IER cases, we also noticed that LLMs mostly failed to reason about loop conditions correctly. Without understanding loop conditions and understanding the number of iterations, it is impossible to reason about the execution correctly. Furthermore, we found cases where, although the model could reason about loop conditions and the number of iterations, it lost track of data flow in the loop, thereby predicting output incorrectly.
(a) Java
(b) Python
Figure 7: Performances of LLMs on code with different types of outputs. Type Match(TM) means the ratio of predicted output that is in the same type as the expected output (ground truth). Value Match(VM) means the CRR on the code with a certain type of output.
In the code snippet below (p03059 from CodeNet (Java)), the loop condition depends on a constant variable (c) and variable (time), whose value changes inside the loop.
Input: 3 5 7, Expected Output:10
int a, b, c = sc.nextInt();
int amount = 0, time = a;
for(int i = 2; time <= c + 0.5; i++){
    amount += b;
    time = a * i;
println(amount);
All models failed in IER of this relatively simple program. Looking at the CoT analysis, GPT-3.5 correctly understands the loop condition and can reason about all the variable values in the first and second iterations. However, it stays in the loop for one more. StarCoder cannot correctly reason about loop condition and iterates the code block six times. MagicCoder CoT only explains the code (correct explanation) without communicating how it reasons about variables.
GPT-3.5, Predicted Output:15
Analysis: Enters for loop and iterates from 2 onwards
until the value of time exceeds c + 0.5. In first
iteration, amount becomes 5 and time becomes 6. In
second iteration, amount becomes 10 and time becomes
9. In third iteration, amount becomes 15 and time
becomes 12. 12 is greater than 7+0.5, loop exits@.
StarCoder, Predicted Output:105
Analysis: It enters a for loop and it iterates from 2
to 7 to check each integer i in this range. In each
iteration, the value of i is passed to variable time.
The condition is@ met when i is 2,3,4,5,6 and 7.
MagicCoder, Predicted Output:15
Enters for loop and iterates from 2 to infinity. Inside
the loop, it adds the value of b to variable amount.
Then it calculates new value of time by multiplying a
with current iteration i. The loop continues until time
is greater than c+0.5. After the loop finishes, print
the final value of amount as output, which is 15.
The loop condition is simple in the code snippet below (p02994 from CodeNet (Java)). Consequently, most models correctly reasoned about the number of iterations.
Input: 5 2, Expected Output: 18
int n = scanner.nextInt();
int l = scanner.nextInt();
List<Integer> diff = new ArrayList<Integer>();
for(int i = 1; i < n+1; i++) {
    int value = l+i-1;
    if(value < 0) {value = value * -1;}
    diff.add(value);
removeK = diff.indexOf(Collections.min(diff)) + 1;
println(n*l + n*(n+1)/2 - n - (l+ removeK -1));
Still, some LLMs struggled to reason about code execution inside the loop correctly. CodeLlama can correctly reason about the number of loop iterations but fails to calculate the values in the diff array, resulting in a wrong output. MagicCoder can also correctly reason about loop length, and similar to the previous example, correctly explain what the code does. However, it cannot predict the output correctly (predicted output is 10), preventing us from judging if it understands the code as well as it explains it.
CodeLlama, Predicted Output:24
Analysis: In the given inputs n is 5 and l is 2. The
sequence of numbers from 1 to 5 is [1,2,3,4,5]. The
differences between l and each number in the sequence
are [1,0,1,2,3]. The smallest difference is 0 at index
1. So, the index 1 is removed from the sequence. The
new sequence becomes [1,3,4,5]. The sum of the original
sequence [1,2,3,4,5] and the new sequence [1,3,4,5] is
calculated as (5*2+5*(5+1)/2-5-(2+1-1))=15+15-5-2=24.
Impact of if-else Properties. Our manual investigations reveal that LLMs struggle with conditional constructs for two reasons: (1) confusion about branches and (2) wrong reasoning about conditional predicates. Concerning the former, we observed that LLMs sometimes struggle to understand the meaning of conditional statements. For example, regardless of which branch to take, they always go into both “if” and “else” branches to reason about execution. However, such cases rarely happen, and in most cases, LLMs struggle to analyze the conditional predicates correctly, take the wrong branch, and predict the output incorrectly.
Impact of Math and Logic Operators. Understanding the meaning of math and logic operators is essential to reasoning how inputs evolve into output through execution. The observations from previous analyses also underscore their importance to reason about the program control flow. Our manual investigation of failed IER cases shows the prevalence of arithmetic and logic operators in programs, with the former dominating the number of the latter. LLMs struggled more to reason about bitwise operators (e.g., $>$, $\wedge$, >>, <<, |, and &\&&) compared to arithmetic (e.g., + and -) and comparison operators (e.g., >, <).
The code snippet below is from MBPP (task 311) and contains several arithmetic and logic operators.
Input: 10, Expected Output:14
def set_left_most_unset_bit(n):
  if not (n & (n + 1)): return n
  pos, temp, count = 0, n, 0
  while temp:
    if not (temp & 1): pos = count
    count += 1
    temp>>=1
  return (n | (1 << (pos)))
StarCoder, Predicted Output:10
Analysis: The variable n is initialized to 10. The code
checks if n is a power of 2. If it is, the code returns
n. If it is not, the code calculates the position of
the leftmost unset bit. The code then returns the
result of n | (1 << (pos)).
StarCoder, while failing to predict the output, correctly understands the meaning of most operators. For example, it can reason that the bitwise AND of n and n+1 is used to determine whether integer n is a power of 2. It can also reason that the loop condition checks if the least significant bit of n is set to 0. However, it seemingly cannot understand the meaning of the return statement (setting the least significant bit to 1), hence failing to predict the correct output.
Impact of Output Types. We categorized programs based on the output types and checked (1) if LLMs were able to correctly predict the type of output (Type Match) and (2) if they could correctly reason about the values of output (Value Match). We identified seven types in subject programs, namely Int (e.g., $2$), Decimal(e.g., $2.34$), String (e.g., ”CodeMind”), Binary (e.g., True or False), List (e.g., [1,3,4,7]), and Tuple (Python-specific, e.g., (2,7)). Figure 7 shows the details of these results. In summary, LLMs achieve a high Type Match ($>80\%$), although they struggled to predict the correct value (Value Match). Among different types, it is harder for the models to predict the values of outputs with Tuple/List and Decimal types.
Tuples and Lists consist of multiple items and every single one of them may change during the program execution. As a result, it is unsurprising that models struggle to track the flow of inputs through potentially different execution paths and reason about a complex output as a whole. Additionally, given that manipulation of such types involves API calls, e.g., min(), next(), charAt(), understanding changes requires additional efforts by LLMs.
9 Concluding Remarks
In this paper, we discussed the necessity of code reasoning tasks as an alternative to evaluate LLMs for programming tasks. We introduced CodeMind, a framework that supports several code reasoning tasks, and used CodeMind in a large-scale grounded theory study to evaluate state-of-the-art LLMs for code reasoning. Our results demonstrate that LLMs, in general, understand code constructs and are capable of reasoning about program specification and following how inputs evolve to output through execution. However, their ability is limited as the code becomes more complex, i.e., has more complex control- or data flow, contains non-primitive types, and invokes API calls. We also observe that specification reasoning, which is essential to generate a code from a given program specification, does not mean models can also reason about code execution. We are considering two future directions based on this work. First, we plan to add more code reasoning tasks to CodeMind, e.g., variable reasoning and code optimization reasoning. Furthermore, we want to augment CodeMind with a benchmark that can challenge LLMs’ code reasoning to a greater extent compared to the existing benchmarks.